OpenStudy (anonymous):

square root of 300 divided by the square root of 5

OpenStudy (anonymous):

sqrt 300 is same as sqrt 5 x sqrt60 if u have done surds

OpenStudy (anonymous):

so sqrt 5 cancels and ur left with sqrt 60

OpenStudy (anonymous):

This is like the other one, combine sqrt(300/5) = sqrt (60) now uncombine = sqrt(4*15) = 2*sqrt(15)